Please show ALL of your work!

Please show ALL of your work!
Answer #1

(a) this is asking for the probability that a structural component does not fail, or that the event E1 does not occur. So the probability of E1 occurring is 0.22. The probability of E1 not occurring is
1-0.22=0.78

(b) this is asking the probability that a component fails (event E1), or shows strain (event E2.) Since it is asking the probability of E1 or E2 happening, you add the probabilities together.
P(E1)+P(E2) =0.22+0.43=0.65

(c) this is asking the probability that the events of part (b) of the question do not take place, so (E1) and (E2) don't happen.

=1-(E1)-(E2) = 1-0.22-0.43=0.35